mathslover (mathslover):

find the minimum value of \( 4-5x-4x^2\).

OpenStudy (anonymous):

For any parabola of the form \(ax^2+bx+c=0 \) the vertex will be at \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \).Equation of line of symmetry will be at \(x=-\frac b {2a} \) Maximum and minimum values will be at \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \) and \( \large \left(-\frac b {2a} , -\frac {b^2-4ac}{4a}\right) \) for \(a<0\) and \(a>0\) respectively.

mathslover (mathslover):

@FoolForMath but that does not lead to the correct answer

OpenStudy (anonymous):

i think there is no minimum value

OpenStudy (anonymous):

it's unbounded

OpenStudy (anonymous):

For this problem there will be no minimum values.

mathslover (mathslover):

can we do this like : 4-5x-4x^2 4-4(x^2+5x/4) 4-4(x^2+5x4 +25/64) + 25/16 ?

mathslover (mathslover):

is this the correct method ?

OpenStudy (anonymous):



OpenStudy (anonymous):

just maximum value

mathslover (mathslover):

4-4(x^2+5x4 +25/64) + 25/16 89/16-4((x+5/8)^2) hence it as min. value of 89/16 at x =-5/8 is it right ?

OpenStudy (anonymous):

nope

mathslover (mathslover):

why ?

OpenStudy (anonymous):

THIS FUNCTION HAS NO MINIMUM VALUE

mathslover (mathslover):

http://www.wolframalpha.com/input/?i=4-5x-4x%5E2&t=crmtb01 i am getting right according to this but it may be wrong please have a look

OpenStudy (anonymous):

Lets try to plot this function graphically,

mathslover (mathslover):

k

OpenStudy (anonymous):

ok try x = 10000

OpenStudy (anonymous):

when do you think the minimum value will be achieved?

OpenStudy (anonymous):

why?

mathslover (mathslover):

can u tell m not soo good

OpenStudy (anonymous):

every x you get i will pick one,that that my function will be less than yours

OpenStudy (anonymous):

Yes, for every minimum you find I can find another x which gives a lesser value ...

mathslover (mathslover):

thanks a lot every one